OpenStudy (anonymous):

The surface area of a cube is 100 square centimeters. Estimate to the nearest tenth of a centimeter the side length of the cube using the formula S = 6s2, where S represents surface area and s represents side length. Make sure to show all work and label your answer.

OpenStudy (cwrw238):

100 = 6s^2 solve for s

OpenStudy (anonymous):

6s^2=100 log6s^2=log10^2 log6+logs^2=2log10 2log s=2-log 6 log s=(2-log6)/2 s=4.08

OpenStudy (cwrw238):

no need to use logs s^2 = 100/6 = 16.666 s = sqrt 16.666 = 4.08

OpenStudy (anonymous):

oh true i was doing lots of log and cannot turn it back btw congrad for getting the answer